A gas occupies a certain volume at 27 degree celcius.If it is heated at constant pressure, its volume is exactly doubled at a temperature of
1
Expert's answer
2012-04-03T09:46:54-0400
If pressure is constant then due to Gay-Lussac's law& V/T=const
So V/T=2V/x => x=2V*T/V=2*T=2*27=54 degrees celcius